# Shuffle Products, Degenerate Affine Hecke Algebras, and Quantum Toda Lattice

### v2

@@ Appendix B Dual representation
$\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega^{*}]/(A(\Omega^{*}-\hbar))\xrightarrow{\sim}\mathrm{U}_{\hbar}(\mathfrak{gl}^{R}_{N})\otimes(\mathbf{C}^{N})^{*}/\!\!/_{\psi}\,N_{+}^{R}\xrightarrow{\sim}H^{\bullet}_{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}(\mathrm{Gr}^{\omega_{1}}),$
that coincides with the Bezrukavnikov-Finkelberg isomorphism (B.9).
@@
$\mathrm{Gr}^{\omega_{1}}\cong\mathbb{P}[(\mathbf{C}^{N})^{*}],\qquad[L_{-1}\subset L_{0}]\mapsto[(L_{0}/L_{-1})^{*}\subset(L_{0}/zL_{0})^{*}],$
@@
$H^{\bullet}_{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}(\mathrm{Gr}^{\omega_{1}})\cong\mathbf{C}[\hbar][A_{1},\ldots,A_{N}]/(A(c_{1}(\mathcal{Q}))),$
The rest of the proof goes as in Proposition˜7.13 (recall that the right module structure on the target is given by (B.2)). ∎

### v4

@@ Appendix B Dual representation
$\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega^{*}]/(A(\Omega^{*}-\hbar))\xrightarrow{\sim}\mathrm{U}_{\hbar}\bigl(\mathfrak{gl}^{R}_{N}\bigr)\otimes\bigl(\mathbf{C}^{N}\bigr)^{*}/\!\!/_{\psi}\,N_{+}^{R}\xrightarrow{\sim}H^{\bullet}_{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}(\mathrm{Gr}^{\omega_{1}})$
that coincides with the Bezrukavnikov–Finkelberg isomorphism (B.8).
@@
$\mathrm{Gr}^{\omega_{1}}\cong\mathbb{P}\bigl[\bigl(\mathbf{C}^{N}\bigr)^{*}\bigr],\qquad[L_{-1}\subset L_{0}]\mapsto[(L_{0}/L_{-1})^{*}\subset(L_{0}/zL_{0})^{*}].$
@@
$H^{\bullet}_{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}(\mathrm{Gr}^{\omega_{1}})\cong\mathbf{C}[\hbar][A_{1},\dots,A_{N}]/(A(c_{1}(\mathcal{Q}))).$
The rest of the proof goes as in Proposition 8.14 (recall that the right module structure on the target is given by (B.2)). ∎
Acknowledgements
The author would like to thank Leonid Rybnikov, Joel Kamnitzer, Michael Finkelberg, Alexis Leroux-Lapierre, and Théo Pinet for helpful discussions and explanations. We would also like to thank the anonymous referees whose comments greatly improved the paper. This work was supported by Fondation Courtois.